A group of the following students got the following score in a test:6,9,12,15,and18 compute the mean of the sample mean

1
Expert's answer
2022-05-18T08:52:42-0400

We have population values 6,9,12,15,18, population size N=5 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{6+9+12+15+18}{5}=12"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{36+9+0+9+36}{5}=18"


"\sigma=\sqrt{\sigma^2}=\sqrt{18}\approx4.24264"

Select a random sample of size 3 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_3=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 6,9,12 & 9 \\
 \hdashline
 2 & 6,9,15 & 10 \\
 \hdashline
 3 & 6,9,18 & 11 \\
 \hdashline
 4 & 6,12,15 & 11 \\
 \hdashline
 5 & 6,12,18 & 12 \\
 \hdashline
 6 & 6,15,18 & 13 \\
 \hdashline
 7 & 9,12,15 & 12 \\
 \hdashline
 8 & 9, 12,18 & 13 \\
 \hdashline
 9 & 9,15,18 & 14 \\
 \hdashline
 10 & 12, 15,18 & 15 \\
 \hdashline
\end{array}"



B.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 9 & 1/10 & 9/10 & 81/10 \\
 \hdashline
 10 & 1/10 & 10/10 & 100/10 \\
 \hdashline
 11 & 2/10 & 22/10 & 242/10 \\
 \hdashline
 12 & 2/10 & 24/10 & 288/10 \\
 \hdashline
 13 & 2/10 & 26/10 & 338/10 \\
 \hdashline
 14 & 1/10 & 14/10 & 196/10 \\
 \hdashline
 15 & 1/10 & 15/10 & 225/10 \\
 \hdashline
\end{array}"



Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{120}{10}=12=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{1470}{10}-(12)^2=3= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{3}\approx1.73205"